OpenStudy (anonymous):

braycub How you solve this algebraically 2/3x+4/x=7/x+1? 16 minutes ago

OpenStudy (anonymous):

To clarify: the equation is\[\frac{ 2 }{ 3x }+\frac{ 4 }{ x }=\frac{ 7 }{ x} + 1\]correct?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

To solve it, then, you would first want to multiply the equation by x to get rid of the x's in denominators. This gives \[\frac{ 2 }{ 3 }+4=7+x\] which can then be solved, giving \[x=\frac{ 2 }{ 3 }+4-7\]

OpenStudy (anonymous):

So that's the ansqer

OpenStudy (anonymous):

answer